### v1

# Discrete optimal transport is a strong audio adversarial attack

Abstract
In this paper, we show that discrete optimal transport (DOT) is an effective black-box adversarial attack against modern audio anti-spoofing countermeasures (CMs). Our attack operates as a post-processing, distribution-alignment step: frame-level WavLM embeddings of generated speech are aligned to an unpaired bona fide pool via entropic OT and a top- barycentric projection, then decoded with a neural vocoder. Evaluated on ASVspoof2019 and ASVspoof5 with AASIST baselines, DOT yields consistently high equal error rate (EER) across datasets and remains competitive after CM fine-tuning, outperforming several conventional attacks in cross-dataset transfer. Ablation analysis highlights the practical impact of vocoder overlap. Results indicate that distribution-level alignment is a powerful and stable attack surface for deployed CMs.
Index Terms— Optimal transport, adversarial attack, ASVspoof
1 Introduction
Vector-embedding voice conversion (VC) with WavLM model [1] began with a simple -nearest neighbors (NN) mapping in embedding space [2]. Subsequent work replaced NN with optimal transport (OT), improving alignment between source and bona fide target distributions and refining conversion quality [3, 4]. In particular, discrete OT (DOT) with barycentric projection was recently applied as a post-processing domain-alignment step on generated audio, yielding scores close to genuine recordings under strong countermeasures (CMs) [4]. OT has likewise been explored on the defense side for distribution comparison in anti-spoofing [5].
A large body of work has shown that automatic speech recognition (ASR) and speaker verification (ASV) pipelines are vulnerable to adversarial examples, including optimization-based attacks like [6] (see also [7, Sec. 7]). In parallel, the ASVspoof challenges established standardized corpora and protocols (e.g., ASVspoof2019 [8]; ASVspoof5 [7]) for training and evaluating spoofing CMs. Among modern CMs, AASIST systems [9] and its modifications leverage spectro-temporal graph attention to capture diverse spoofing artifacts and have become strong baselines for deepfake/anti-spoofing detection. Despite this progress, transferability across datasets and generative models remains an open challenge.
We study whether discrete OT-based voice conversion is a strong black-box attack on state-of-the-art CMs (Figure 1). Our method aligns WavLM frame embeddings to an unpaired bona fide pool via entropic DOT with a cosine cost, applies a top- barycentric map, and reconstructs with a neural vocoder. Unlike gradient attacks, DOT induces a distributional shift toward bona fide regions, enabling cross-dataset transfer without CM internals. Our Contributions:
DOT as an adversarial attack. We formalize discrete OT + barycentric projection VC as a black-box, distribution-alignment attack.
Transfer and fine-tuning robustness. DOT remains strong across datasets ASVspoof2019 and ASVspoof5 and remains competitive after CM fine-tuning, outperforming several conventional attacks.
Ablations and practical factors. We show that vocoder overlap with CM training data modulates attack strength, informing both attack design and CM training.
2 Problem Setup and Threat Model
We consider an audio assistant/ASV pipeline with a spoofing CM trained under ASVspoof protocols [8, 7]. Given waveform , the CM outputs a score ; a threshold yields a binary (bona fide/spoof) decision.
Adversary Goal. Transform a generated/converted input into an attacked sample that (i) is accepted as bona fide by the CM and (ii) preserves intelligibility/naturalness to avoid human suspicion.
Adversary Knowledge and access. We assume a black-box CM: no gradients or internals; score/label access is optional and not required. The adversary may synthesize using any upstream TTS/VC system, and has an unpaired bona fide speech pool from the target domain or a close proxy, reflecting realistic deployments and standard adversarial-audio practice.
Adversary Capabilities (DOT attack). The adversary embeds into frame-level vectors embeds an unpaired bona fide pool , computes an entropic discrete OT coupling matrix under cosine cost , applies a top- barycentric projection to obtain transported embeddings , and reconstructs a waveform via a neural vocoder. The pipeline uses only generic pretrained components and unpaired speech.
Perceptual Constraints. The attack seeks minimal perceptual deviation from while achieving CM evasion.
We follow ASVspoof protocols [8, 7]. Success is demonstrated by higher EER (or higher false-acceptance rate) on than baselines, including under cross-dataset transfer and after CM fine-tuning. We also analyze practical factors — vocoder overlap, top-, and target duration — that modulate attack strength (see Sec 5).
3 Methodology: Discrete OT Attack
3.1 Discrete OT and Barycentric Projection
Let and be two probability spaces. Denote by all joint distributions on the product space with marginals and In discrete case assume there are vectors in and vectors in with probability masses and The joint distribution is represented as a non-negative matrix with and
The goal of optimal transport (OT) is to find the joint distribution known as Kantorovich plan, that minimizes the expected transport cost
subject to the marginal constraints:
Given a solution , a transport map can be defined via the barycentric projection:
This transform can be interpreted as the conditional expectation
3.2 Adversarial Attack
We follow the discrete OT voice-conversion framework of [4]. Let one audio recording be represented by a sequence of embeddings Here, are embeddings obtained with WavLM Large pretrained model [1]. This model encodes every ms of audio into a -dimensional vector embedding, with a hop size of ms.
For every pair of audio recordings from the source and target speakers respectively, we extract their vectorized representations: and with
Since the underlying distributions of speaker embeddings and are unknown, we use empirical distributions: and It was noticed that it is better to have a longer recording for the target speaker; therefore, can contain embeddings from several audio samples.
The goal is to enhance generated audio applying the OT form embeddings to a real sample In other words, we want to define a mapping (transportation map) such that where has a similar distribution to (but it is not necessary an element of ). In the discrete OT approach, we compute the coupling matrix for the given marginal distributions and a cost function. For each we sort the target embeddings in decreasing order of denoting the sorted vectors as The sorted coupling weights along each row (with fixed ) are denoted by
We define the mapping as the baricentric projection of the OT map over top- vectors,
Note that this formula coincides with (3) only when . We use as it was shown in [4], the quality of conversion does not change much for but it degradates for large After the transformation we convert the predicted embeddings back into waveform using HiFi-GAN vocoder (see Sec. 4).
The intuition behind this attack is simple: the approximate map shifts the empirical distribution of generated speech toward the target (real) distribution. Since countermeasure systems are trained to reject synthetic distributions and accept real speech, moving the generated distribution closer to the real one can reduce detection performance — i.e., produce a strong, dataset-agnostic adversarial effect.
3.3 Attack Variants
As the set we used generated audios from ASVspoof2019 dataset (validation part). As the target space we investigated two options: hemi-anechoic chamber recordings and diverse recording conditions. As the result, we considered the following attacks.
OT1:
As the target space we used bonafide recordings from the ASVspoof2019 dataset built on the VCTK [10].
OT2:
As the target space we used the LibriSpeech train-clean-100 dataset [11], because ASVspoof5 used LibriVox data that contains Libri Speech. We select the first 40 speakers (ordered by speaker ID) and, for each, extract 10 random utterances and sort them by duration.
It was noticed in previous research (see, e.g., [4]) that the quality of the voice conversion depends on the length of the target data. Using LibtiSpeech dataset was could concatenate several utterances for the same speaker. Evaluation showed that the difference between using VCTK or LibriSpeech is not large (see Table 1), we decided to use OT2 as the main attack for further analysis.
4 Experimental Setup
Datasets. We use three public corpora. For constructing the bona fide target pool in our DOT attack (OT2), we draw from the LibriSpeech Clean, available on Kaggle [12]. For countermeasure (CM) evaluation and cross-dataset transfer, we use ASVspoof2019 accessible on [13], and ASVspoof5 [7]. Unless otherwise stated, we use the official train/dev/eval partitions of each benchmark.
Embeddings. We extract frame-level embeddings using WavLM Large, taking embeddings from the sixth transformer layer accessed through KNN-VC framework [3].
Optimal transport. Discrete OT is solved with entropic regularization via the Sinkhorn algorithm from the POT library [14]. The cosine cost is used; the regularization parameter is set to Top- barycentric projection produces transported embeddings with
Vocoder. We reconstruct waveforms from transported embeddings with HiFi-GAN, using the implementation provided with the KNN-VC framework [3].
Countermeasure. For evaluation, we adopt the official AASIST implementation [9] and its pretrained variants.
Evaluation Metric 1: EER We report the equal error rate (EER), i.e., the operating point where the false acceptance rate equals the false rejection rate (see, e.g., [15, Sec. 13]).
5 Analysis and Evaluation
For evaluation, we used the AASIST model [9], pretrained on the ASVspoof2019 dataset (denoted as AASIST2019), and on ASVspoof5 (denoted as AASIST5).
Table 1 reports equal error rates (EER) for the generation algorithm A18 from ASVspoof2019 and attack A18 in ASVspoof5, which we denote by A185, as well as for the two attacks introduced in the previous section. Algorithms A18 and A185 were selected because they exhibit the highest EER within their respective datasets.
| Attack | AASIST2019 | AASIST | AASIST5 | AASIST |
| A18 | 2.614 | 3.141 | 77.735 | 44.951 |
| A185 | 0.435 | 1.443 | 57.933 | 2.730 |
| OT1 | 11.111 | - | 7.268 | - |
| OT2 | 7.925 | 0.216 | 11.180 | 12.586 |
Column AASIST2019 shows the EER on the validation subset of ASVspoof2019, evaluated with AASIST2019 (i.e., using bonafide data from the ASVspoof2019 validation set).
Column AASIST5 presents the EER computed with AASIST5. A notable observation emerges: the strongest attack proposed in ASVspoof5 is detected robustly by the model pretrained on ASVspoof2019. Conversely, most methods from ASVspoof2019 achieve higher EERs than the novel attacks in ASVspoof5 (see also columns (2019) and (5) in Table 2). This asymmetry reflects adversarial transferability across models/datasets [18, 19, 20], observed previously in speech systems as well [21, 22].
Column AASIST shows results after fine-tuning AASIST2019 by including OT2 examples in the training set of ASVspoof2019.
The final column reports results after fine-tuning AASIST5 using OT2 and A185 data (see Sec. 5.1 for details).
Table 2 extends this analysis by including attacks A07–A19 (ASVspoof2019) and A175–A315 (ASVspoof5). Columns (2019) and (5) represent baseline evaluation with AASIST2019 and AASIST5, while columns (2019OT) and (5OT) show results after applying optimal transport to A07–A19 data.
5.1 Fine-Tuning
Since all generation methods and attacks from ASVspoof5 resulted in very low EER when detected by AASIST2019, we fine-tuned AASIST2019 only with OT2 data. These examples were obtained by applying optimal transport to generated audio in the ASVspoof2019 training set.
Comparing columns (2019OT) and (2019) in Table 2 (equivalently, columns AASIST2019 and AASIST in Table 1), we see that OT2 attacks are easily detected after fine-tuning. The EER for A185 remained largely unchanged.
For AASIST5, which suffers from a strong A18 attack (and also from A19, though here we focus on adversarial attacks), we fine-tuned using a subset of A185 data from the ASVspoof5 evaluation set (12,000 of 27,000 recordings, with the remainder reserved for evaluation) and included OT2 training data used for fine-tuning AASIST2019.
The comparison between AASIST5 and AASIST (Table 1) shows that A185 becomes well detected after fine-tuning. However, the OT2 attack still maintains a high EER.
5.2 The Role of the Vocoder
The relatively low EER for OT2 in AASIST5 (Table 1) compared with other attacks (see also Table 2, column (5), rows A175–A315) may be explained by vocoder overlap. Most of the ASVspoof5 training data are generated with the HiFi-GAN vocoder, which is also used in our attack. This likely explains the elevated EER for methods employing different vocoders, particularly the extremely high EER observed for ASVspoof2019 methods evaluated with AASIST5 (Table 2, column (5), rows A07–A19).
| Attack | 2019 | 2019 | 5 | 5 | ||
| A07 | 0.52 | 0.51 | 0.12 | 33.28 | 4.39 | 10.05 |
| A08 | 0.42 | 3.70 | 0.11 | 19.90 | 4.23 | 9.49 |
| A09 | 0.00 | 0.69 | 0.05 | 7.79 | 2.92 | 7.44 |
| A10 | 0.86 | 0.73 | 0.13 | 39.42 | 6.39 | 14.30 |
| A11 | 0.18 | 0.71 | 0.13 | 5.65 | 5.40 | 13.71 |
| A12 | 0.78 | 0.73 | 0.09 | 55.41 | 7.40 | 13.04 |
| A13 | 0.15 | 0.51 | 0.06 | 65.51 | 6.38 | 15.07 |
| A14 | 0.15 | 0.81 | 0.05 | 15.34 | 4.34 | 8.15 |
| A15 | 0.55 | 0.77 | 0.05 | 7.30 | 3.94 | 8.88 |
| A16 | 0.65 | 1.98 | 0.11 | 71.97 | 11.37 | 17.15 |
| A17 | 1.26 | 13.34 | 0.37 | 73.26 | 17.69 | 12.36 |
| A18 | 2.61 | 12.26 | 0.27 | 77.73 | 24.19 | 14.74 |
| A19 | 0.65 | 12.98 | 0.27 | 72.57 | 19.02 | 14.07 |
| A175 | 1.36 | - | - | 11.43 | - | - |
| A185 | 0.43 | - | - | 51.92 | - | - |
| A195 | 0.18 | - | - | 57.93 | - | - |
| A205 | 0.11 | - | - | 49.78 | - | - |
| A215 | 0.72 | - | - | 13.28 | - | - |
| A225 | 1.24 | - | - | 14.07 | - | - |
| A235 | 0.15 | - | - | 28.81 | - | - |
| A245 | 1.39 | - | - | 10.69 | - | - |
| A255 | 0.35 | - | - | 22.29 | - | - |
| A265 | 1.86 | - | - | 27.41 | - | - |
| A275 | 0.18 | - | - | 24.10 | - | - |
| A285 | 0.94 | - | - | 23.57 | - | - |
| A295 | 0.58 | - | - | 6.83 | - | - |
| A305 | 0.25 | - | - | 39.89 | - | - |
| A315 | 0.18 | - | - | 26.53 | - | - |
5.3 The Optimal Transport Property
To further illustrate the effect of optimal transport, we applied t-SNE to VGGish embeddings.
Figure 2 shows embeddings from LibriSpeech bonafide data, ASVspoof2019 generated data (A01–A06), and their OT-transformed versions. The transformed embeddings (Converted) align closely with the LibriSpeech target distribution (Bonafide).
Figure 3 shows embeddings from ASVspoof2019 bonafide data, ASVspoof2019 generated data, and their OT-transformed counterparts. Here, the discrepancy between ASVspoof2019 bonafide data and LibriSpeech becomes evident. This explains why the Fréchet Audio Distance (FAD) between bonafide training data and OT2 is relatively large (Table 3), and why AASIST2019 readily detects OT2 after fine-tuning, despite its high pre-finetuning EER.
| BF Dataset | BF–Spoof | BF–OT | Spoof–OT |
| LibriSpeech | 4.742 | 0.508 | 3.665 |
| ASVspoof2019 | 1.289 | 3.402 | 3.665 |
6 Acknowledgment
The first author thanks You (Neil) Zhang for pointing to the adversarial attacks in ASVspoof5 dataset and bringing to my attention paper [5].