### v2

# Selitskiy Shahriyar Prakasan Discrete Optimal Transport is a Strong Audio Adversarial Attack

Abstract
In this paper, we introduce the discrete optimal transport voice conversion (DOT-VC) method. Comparison with kNN-VC, SinkVC, and Gaussian optimal transport (MKL) demonstrates stronger domain adaptation abilities of our method. We use the probabilistic nature of optimal transport (OT) and show that DOT-VC is an effective black-box adversarial attack against modern audio anti-spoofing countermeasures (CMs). Our attack operates as a post-processing, distribution-alignment step: frame-level WavLM embeddings of generated speech are aligned to an unpaired bona fide pool via entropic OT and a top- barycentric projection, then decoded with a neural vocoder. Ablation analysis indicates that distribution-level alignment is a powerful and stable attack for deployed CMs.
keywords:
optimal transport, adversarial attack, ASVspoof1 Introduction
1.1 Voice conversion (VC) with WavLM
Vector-embedding-based voice conversion (VC) using the WavLM model [wavlm] was initially introduced via a simple -nearest neighbors (NN) mapping in the embedding space [VCNN]. Let denote the sequence of embeddings extracted from a source speaker recording, where are obtained using the pretrained WavLM Large model. WavLM encodes audio using a window length of ms and a hop size of ms. Similarly, let denote the embeddings of a target speaker utterance.
The goal of NN-VC is to convert into where each is obtained by performing NN regression of over the target embeddings , using cosine distance They set
Subsequent work [VCOT] proposed replacing distance-based NN weights with transport probabilities learned via optimal transport (OT). Instead of averaging neighbors, each source embedding is mapped to an average of four target embeddings corresponding to top-4 This method, referred to as SinkVC in [VCOT], derives its name from the Sinkhorn algorithm commonly used to solve discrete optimal transport (DOT) problems.
In addition, the authors of [VCOT] train a neural network on pairs and refer to this approach as flow matching VC (FMVC). We find this terminology potentially misleading: in the original flow matching framework [flowm], one of the distributions is assumed to be standard Gaussian, and the method does not explicitly learn an OT map. In contrast, their setup relies on precomputed OT pairs, which allows one to learn geodesics (see, e.g., [Anbrosio, Theorem 11.2.10]).
To reduce computational cost, [got] proposed an approximation based on continuous OT. Specifically, the authors estimate means and covariance matrices of the source and target embeddings, respectively, and apply the closed-form Gaussian OT map
This corresponds to optimal transport between two Gaussian distributions. In our experiments, this approach performs poorly, which may explain the introduction of a variance-stratified scheme in [got].
In that scheme, the standard deviations of individual embedding components are computed, and the embedding dimensions are sorted in decreasing order of . The reordered embeddings and are then partitioned into blocks of equal size , and continuous OT is applied independently to each block. The authors consider and refer to this method as factorized MKL (Monge—Kantorovich Linear). While this approach significantly improves computational efficiency and can be applied to short utterances, treating embedding components as uncorrelated across blocks leads to a noticeable loss of target speaker identity (see Table 1).
In contrast to the above methods, we propose using a truncated barycentric projection of discrete OT. The barycentric projection is defined as
By retaining only the largest transport normalized weights , defined in Sec. 2.1, we obtain the proposed DOT-VC method. We use ; as a preliminary preprint [DVCOT] shows, the conversion quality remains stable for and degrades for larger values of . Finally, our ablation study demonstrates that DOT-VC more accurately approximates the target speaker distribution than NN-VC or MKL methods (see Table 1).
1.2 Automatic speaker verification (ASV) systems
A substantial body of prior work has demonstrated that automatic speech recognition (ASR) and automatic speaker verification (ASV) pipelines are vulnerable to adversarial examples, including optimization-based attacks such as [malafide] (e.g., in Table 4, see [wang2024asvspoof, Sec. 7]). In parallel, the ASVspoof challenges have established standardized datasets and evaluation protocols (e.g., ASVspoof2019 [wang2020asvspoof], ASVspoof5 [wang2024asvspoof]) for training and benchmarking spoofing countermeasure (CM) systems.
Among modern CMs, AASIST [Jung2021AASIST] and its variants have emerged as strong baselines for deepfake and anti-spoofing detection. Despite significant progress, generalization and transferability across datasets, recording conditions, and generative models remain open challenges for current CM systems.
We investigate whether DOT-based voice conversion constitutes a strong black-box attack against SOTA CMs (Figure 1). Our method aligns WavLM frame-level embeddings of a generated utterance to an unpaired bona fide embedding pool using entropic DOT with a cosine cost, applies a top- barycentric projection, and reconstructs audio using a neural vocoder. Unlike gradient-based attacks, DOT induces a distributional shift toward bona fide regions of the embedding space, enabling effective cross-dataset transfer without access to CM internals.
1.3 Problem setup and threat model
We consider an audio assistant or ASV pipeline equipped with a spoofing CM. Given an input waveform , the CM outputs a score , which is compared against a threshold to yield a binary decision (bona fide vs. spoof).
Adversary goal. Given a generated utterance and a real speech sample , the adversary aims to transform into an attacked sample such that (i) is accepted as bona fide by the CM, and (ii) intelligibility and naturalness are preserved (and, when represents a different speaker, the acoustic characteristics of are approximated).
Adversary knowledge and access. We assume a black-box CM: the adversary has no access to model internals or gradients, and score or label access is optional and not required. The adversary may generate using any upstream TTS or VC system and possesses an unpaired pool of bona fide speech from the target domain or a close proxy. This assumption reflects realistic deployment scenarios and aligns with standard adversarial-audio threat models.
Adversary capabilities (DOT attack). The adversary embeds into frame-level representations and embeds an unpaired bona fide pool . An entropic discrete OT problem is solved to obtain a coupling matrix under the cosine cost A top- barycentric projection is then applied to produce transported embeddings , which are converted back to a waveform using a neural vocoder. The pipeline relies solely on generic pretrained components and unpaired speech data.
Perceptual constraints. The attack aims to minimize perceptual deviation from while preserving linguistic content from and achieving CM evasion. We do not impose explicit perturbation norms (e.g., constraints), as the proposed method operates via distributional alignment rather than pointwise perturbations.
We follow ASVspoof evaluation protocols [wang2020asvspoof, wang2024asvspoof]. Attack success is demonstrated by increased equal error rate (EER) or false acceptance rate on attacked samples compared to baselines, including under cross-dataset transfer and after CM fine-tuning. We further analyze practical factors — vocoder overlap. For the influence of top-, and target utterance duration see [DVCOT].
1.4 Our contributions
A novel voice conversion method. We introduce DOT-VC, a voice conversion approach based on truncated barycentric projection of discrete optimal transport.
DOT as an adversarial attack. We formalize DOT-VC as a black-box, distribution-alignment attack against spoofing countermeasures.
Transferability and robustness to fine-tuning. We show that DOT-based attacks generalize across ASVspoof2019 and ASVspoof5 and remain effective after CM fine-tuning, outperforming several conventional attacks.
Ablations and practical insights. We compare WavLM-based VC methods in terms of speech quality (WER, MOS) and domain adaptation metrics (FAD). We compare our attack with attacks in ASVspoof5 dataset. Based on the analysis, we hypothesize that vocoder overlap with CM training data significantly modulates attack strength, with implications for both attack design and CM training.
The code and audio samples can be found on https://anton-selitskiy.github.io/DOTattack.
2 Methodology: Discrete OT Attack
2.1 Discrete OT and barycentric projection
Assume there are vectors in and vectors in with probability masses and The joint distribution is represented as a non-negative matrix with and The goal of optimal transport (OT) is to find the joint distribution known as Kantorovich plan, that minimizes the expected transport cost
subject to the marginal constraints:
Given a solution , a transport map can be defined via the barycentric projection (1), where This transform can be interpreted as the conditional expectation
Since the underlying distributions of speaker embeddings are unknown, we use empirical distributions: and For each we sort the target embeddings in decreasing order of denoting the sorted vectors as The sorted coupling weights along each row (with fixed ) are denoted by We define the DOT mapping as the baricentric projection of the OT map over top- vectors,
| NN | SinkVC | DOT | MKL-2 | MKL-256 | |
| FAD | 0.808 | 0.790 | 0.775 | 1.173 | 1.389 |
The intuition behind this attack is simple: the approximate map shifts the empirical distribution of generated speech toward the target (real speech) distribution. Since CMs are trained to reject synthetic distributions and accept real speech, moving the generated distribution closer to the real one can reduce detection performance, i.e., produce a strong, dataset- and model-agnostic adversarial effect.
In Table 2, we report WER and MOS before (column one) and after conversion in setup [DVCOT, Sec. 4.1] to demonstrate that the attack preserves speech intelligibility (for 100 random audio samples).
| Libri | NN | DOT | MKL-2 | MKL-256 |
2.2 Adversarial attack variants
As the set we used generated audios from ASVspoof2019 dataset (validation part). As the target space we investigated two options: hemi-anechoic chamber recordings and diverse recording conditions. As the result, we considered the following attacks.
OT1: As the target space we used bonafide recordings from the ASVspoof2019 dataset built on the VCTK [vctk].
OT2: As the target space we used the LibriSpeech train-clean-100 dataset [panayotov2015librispeech], because ASVspoof5 used LibriVox data that contains LibriSpeech. We select the first 40 speakers (ordered by speaker ID) and, for each, extract 10 random utterances and sort them by duration.
It was noticed in previous research (see, e.g., [DVCOT]) that the quality of the voice conversion depends on the length of the target data. Using LibriSpeech dataset we could concatenate several utterances for the same speaker. Evaluation showed that the difference between using VCTK or LibriSpeech is not large (see Table 3), we decided to use OT2 as the main attack for further analysis.
3 Experimental Setup
Datasets. We use three public corpora. For constructing the bona fide target pool in our DOT attack (OT2), we draw from the LibriSpeech Clean, available on Kaggle [kaggleLS]. For countermeasure (CM) evaluation and cross-dataset transfer, we use ASVspoof2019 accessible on [kaggleAS] and ASVspoof5 [wang2024asvspoof]. Unless otherwise stated, we use the official train/dev/eval partitions of each benchmark.
Embeddings. We extract frame-level embeddings using WavLM Large, taking embeddings from the sixth transformer layer accessed through NN-VC framework [VCOT].
Optimal transport. Discrete OT is solved with entropic regularization via the Sinkhorn algorithm from the POT library [pot]. The default regularization parameter is used.
Vocoder. We reconstruct waveforms from transported embeddings with HiFi-GAN, using the implementation provided with the NN-VC framework [VCOT].
Countermeasure. For evaluation, we adopt the official AASIST implementation [Jung2021AASIST] and its pretrained variants.
Evaluation metric 1: EER We report the equal error rate (EER), i.e., the operating point where the false acceptance rate equals the false rejection rate (see, e.g., [eer, Sec. 13]).
Evaluation metric 2: distributional similarity. To quantify distributional alignment effects of DOT independently of a specific CM, we compute the Fréchet Audio Distance (FAD) [fadtk]. We use torchvggish (v0.2) for VGGish embeddings [vgg], see details in [DVCOT].
Evaluation metrics 3: speech quality. To compute word error rate (WER), we used the Whisper base ASR model. MOS was automatically computed using the code provided with the UTMOSv2 paper [utmosv2].
4 Analysis and Evaluation
For evaluation, we used the AASIST model [Jung2021AASIST], pretrained on the ASVspoof2019 dataset (denoted as AASIST2019), and on ASVspoof5 (denoted as AASIST5).
Table 3 reports equal error rates (EER) for the generation algorithm A18 from ASVspoof2019 and attack A18 in ASVspoof5, which we denote by A185, as well as for the two attacks introduced in the previous section. Algorithms A18 and A185 were selected because they exhibit the highest EER within their respective datasets.
| Attack | AASIST2019 | AASIST | AASIST5 | AASIST |
| A18 | 2.614 | 3.141 | 77.735 | 44.951 |
| A185 | 0.435 | 1.443 | 57.933 | 2.730 |
| OT1 | 11.111 | - | 7.268 | - |
| OT2 | 7.925 | 0.216 | 11.180 | 12.586 |
Column AASIST2019 shows the EER on the validation subset of ASVspoof2019, evaluated with AASIST2019 (i.e., using bonafide data from the ASVspoof2019 validation set).
Column AASIST5 presents the EER computed with AASIST5. A notable observation emerges: the strongest attack proposed in ASVspoof5 is detected robustly by the model pretrained on ASVspoof2019. Conversely, most methods from ASVspoof2019 achieve higher EERs than the novel attacks in ASVspoof5 (see also columns (2019) and (5) in Table 4). This asymmetry reflects adversarial transferability across models/datasets [szegedy2014intriguing, goodfellow2015explaining, papernot2016transferability], observed previously in speech systems as well [alzantot2018did, asv].
Column AASIST shows results after fine-tuning AASIST2019 by including OT2 examples in the training set of ASVspoof2019.
The final column reports results after fine-tuning AASIST5 using OT2 and A185 data (see Sec. 4.1 for details).
Table 4 extends this analysis by including attacks A07–A19 (ASVspoof2019) and A175–A315 (ASVspoof5). Columns (2019) and (5) represent baseline evaluation with AASIST2019 and AASIST5, while columns (2019OT) and (5OT) show results after applying optimal transport to A07–A19 data.
4.1 Fine-tuning
Since all generation methods and attacks from ASVspoof5 resulted in very low EER when detected by AASIST2019, we fine-tuned AASIST2019 only with OT2 data. Examples were obtained by applying optimal transport to generated audio in the ASVspoof2019 training set.
Comparing columns (2019OT) and (2019) in Table 4 (equivalently, columns AASIST2019 and AASIST in Table 3), we see that OT2 attacks are easily detected after fine-tuning. The EER for A185 remained largely unchanged.
For AASIST5, which suffers from a strong A18 attack (and also from A19, though here we focus on adversarial attacks), we fine-tuned using a subset of A185 data from the ASVspoof5 evaluation set (12,000 of 27,000 recordings to keep the proportion of classes, with the remainder reserved for evaluation) and included OT2 training data used for fine-tuning AASIST2019.
The comparison between AASIST5 and AASIST (Table 3) shows that A185 becomes well detected after fine-tuning. However, the OT2 attack still maintains a high EER.
4.2 The role of the vocoder
The relatively low EER for OT2 in AASIST5 (Table 3) compared with other attacks (see also Table 4, column (5), rows A175–A315) may be explained by vocoder overlap. Most of the ASVspoof5 training data are generated with the HiFi-GAN vocoder, which is also used in our attack. This likely explains the elevated EER for methods employing different vocoders, particularly the extremely high EER observed for ASVspoof2019 methods evaluated with AASIST5 (Table 4, column (5), rows A07–A19).
| Attack | 2019 | 2019 | 5 | 5 | ||
| A07 | 0.52 | 0.51 | 0.12 | 33.28 | 4.39 | 10.05 |
| A08 | 0.42 | 3.70 | 0.11 | 19.90 | 4.23 | 9.49 |
| A09 | 0.00 | 0.69 | 0.05 | 7.79 | 2.92 | 7.44 |
| A10 | 0.86 | 0.73 | 0.13 | 39.42 | 6.39 | 14.30 |
| A11 | 0.18 | 0.71 | 0.13 | 5.65 | 5.40 | 13.71 |
| A12 | 0.78 | 0.73 | 0.09 | 55.41 | 7.40 | 13.04 |
| A13 | 0.15 | 0.51 | 0.06 | 65.51 | 6.38 | 15.07 |
| A14 | 0.15 | 0.81 | 0.05 | 15.34 | 4.34 | 8.15 |
| A15 | 0.55 | 0.77 | 0.05 | 7.30 | 3.94 | 8.88 |
| A16 | 0.65 | 1.98 | 0.11 | 71.97 | 11.37 | 17.15 |
| A17 | 1.26 | 13.34 | 0.37 | 73.26 | 17.69 | 12.36 |
| A18 | 2.61 | 12.26 | 0.27 | 77.73 | 24.19 | 14.74 |
| A19 | 0.65 | 12.98 | 0.27 | 72.57 | 19.02 | 14.07 |
| A175 | 1.36 | - | - | 11.43 | - | - |
| A185 | 0.43 | - | - | 51.92 | - | - |
| A195 | 0.18 | - | - | 57.93 | - | - |
| A205 | 0.11 | - | - | 49.78 | - | - |
| A215 | 0.72 | - | - | 13.28 | - | - |
| A225 | 1.24 | - | - | 14.07 | - | - |
| A235 | 0.15 | - | - | 28.81 | - | - |
| A245 | 1.39 | - | - | 10.69 | - | - |
| A255 | 0.35 | - | - | 22.29 | - | - |
| A265 | 1.86 | - | - | 27.41 | - | - |
| A275 | 0.18 | - | - | 24.10 | - | - |
| A285 | 0.94 | - | - | 23.57 | - | - |
| A295 | 0.58 | - | - | 6.83 | - | - |
| A305 | 0.25 | - | - | 39.89 | - | - |
| A315 | 0.18 | - | - | 26.53 | - | - |
4.3 The optimal transport property
To further illustrate the effect of optimal transport, we applied t-SNE to VGGish embeddings.
Figure 2 shows embeddings from LibriSpeech bonafide data, ASVspoof2019 generated data (A01–A06), and their OT-transformed versions. The transformed embeddings (Converted) align closely with the LibriSpeech target distribution (Bonafide).
Figure 3 shows embeddings from ASVspoof2019 bonafide data, ASVspoof2019 generated data, and their OT-transformed counterparts. Here, the discrepancy between ASVspoof2019 bonafide data and LibriSpeech becomes evident. This explains why the Fréchet Audio Distance (FAD) between bonafide training data and OT2 is relatively large (Table 5), and why AASIST2019 readily detects OT2 after fine-tuning, despite its high pre-finetuning EER.
| BF Dataset | BF–Spoof | BF–OT | Spoof–OT |
| LibriSpeech | 4.742 | 0.508 | 3.665 |
| ASVspoof2019 | 1.289 | 3.402 | 3.665 |
5 Acknowledgment
The first author thanks You (Neil) Zhang for pointing out the adversarial attacks in ASVspoof5 dataset [wang2024asvspoof]. The first author also thanks Arip Asadulaev and Rostislav Korst for clarifications regarding their work [VCOT].
Submitted to Inrterspeech 2026, under review.